# On De Giorgi’s lemma for variational interpolants

### v2

@@ 3.2 De Giorgi’s estimate and identity in the Banach setup
${\mathcal{E}}(u^{o})-{\mathcal{E}}(\widetilde{u}_{\sigma})-\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}\geq\int_{0}^{\sigma}{\mathcal{R}}^{*}\big{(}{-}\widetilde{\xi}_{s}\big{)}\;\!\mathrm{d}s\qquad\text{with }\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ (3.15)
holds for $\sigma\in(0,1]$, while for $\sigma>1$ it only holds for selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ with
$\int_{0}^{\sigma}{\mathcal{R}}^{*}\big{(}{-}\widetilde{\xi}_{s}\big{)}\;\!\mathrm{d}s=\int_{0}^{1}\frac{1}{2}\;\!\mathrm{d}s+\int_{1}^{\sigma}\frac{1}{2}\big{|}{-}\widetilde{\xi}_{s}\big{|}^{2}\;\!\mathrm{d}s\leq 1-\frac{1}{2\sigma}\quad\Leftrightarrow\quad\int_{1}^{\sigma}\big{|}\widetilde{\xi}_{s}\big{|}^{2}\;\!\mathrm{d}s\leq 1-\frac{1}{\sigma}.$
Since $\mathcal{S}_{\mathcal{R}}(\widetilde{u}_{s})=0$ for $s\in[1,\sigma]$, (3.15) is certainly satisfied if we replace the integrand ${\mathcal{R}}^{*}\big{(}{-}\widetilde{\xi}_{s}\big{)}$ by $\mathcal{S}_{\mathcal{R}}(\widetilde{u}_{s})$. In turn, recalling that $\sigma\geq 1$ we have $\partial{\mathcal{E}}(\widetilde{u}_{\sigma})=[0,1]$, a wrong choice of $\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ can violate (3.15).
@@
$0\in\widetilde{\xi}_{\sigma}+\partial{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}\quad\text{with }\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma}).$
@@
$\widetilde{\xi}_{\sigma}=-\frac{1}{\sigma}(\widetilde{u}_{\sigma}-u^{o})=\begin{cases}1&\text{for }\sigma\in(0,1]\,,\\ 1/\sigma&\text{for }\sigma\geq 1\,.\end{cases}$
@@
for $\text{for }\sigma\in(0,1]:$ $\quad\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})=\{1\}\quad\text{and}\quad{\mathcal{S}}_{\mathcal{R}}(\widetilde{u}_{\sigma})={\mathcal{C}}_{\mathcal{R}}(u^{o},\sigma;\widetilde{u}_{\sigma})=\frac{1}{2},$
for $\text{for }\sigma>1:$ $\quad\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})=\big{\{}\frac{1}{\sigma}\big{\}}\quad\text{and}\quad 0={\mathcal{S}}_{\mathcal{R}}(\widetilde{u}_{\sigma})\lneqq{\mathcal{C}}_{\mathcal{R}}(u^{o},\sigma;\widetilde{u}_{\sigma})=\frac{1}{2\sigma^{2}}\>.$
@@
$\widetilde{\xi}_{\sigma}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})\qquad\text{for a.a.}\ \sigma>0\,.$
Thus, the ‘good selections’ of $\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ for the purpose of De Giorgi’s identity is prescribed by the Euler-Lagrange equation (3.6).
@@ 3.3 Radially differentiable potentials
This section revolves around a structural property for dissipation potentials that will be crucial for obtaining De Giorgi’s identity.
@@
We say that a dissipation potential ${\mathcal{R}}:X\to[0,\infty)$ is radially differentiable if
$\forall\,v\in X:\quad(0,\infty)\ni\lambda\mapsto f(v;\lambda):={\mathcal{R}}(\lambda v)\text{ is differentiable at }\lambda=1.$ (3.16)
@@
We are now going to show that radial differentiability is equivalent to another structural property that was used in [MRS13a] to prove De Giorgi’s estimate (3.13).
@@
This ensures that the (a priori multivalued) mapping
[…]
@@
Let us now address the proof of the equivalence of (3.17) and (3.16): We start by observing that for every $v\in X$ the mapping $f(v;\cdot)$ is convex. Its subdifferential in the sense of convex analysis $\partial f(v;\cdot):(0,\infty)\rightrightarrows{\mathbb{R}}$ is given by
$\partial f(v;\lambda)=[f^{\prime}_{-}(v;\lambda),f^{\prime}_{+}(v;\lambda)]\qquad\text{for all }\lambda>0,\ v\in X\,,$ (3.19a)
@@
$f^{\prime}_{+}(v;\lambda):=\lim_{h\to 0^{+}}\frac{1}{h}\big{(}f(v;\lambda{+}h)-f(v;\lambda)\big{)},$ (3.19b)
$f^{\prime}_{-}(v;\lambda):=\lim_{h\to 0^{+}}\frac{1}{h}\big{(}f(v;\lambda)-f(v;\lambda{-}h)\big{)}\,,$
(observe that the above limits exist by convexity of $f(v;\cdot)$).
@@
For every $v\in X$ and $\lambda>0$
@@
Proof. It can be easily checked that, for any $v\in X$ and $\lambda>0$, there holds
@@
Proof of Proposition 3.8: The mapping $f(v;\cdot)$ from (3.16) is differentiable at $\lambda=1$ if and only if $f^{\prime}_{-}(v;1)=f^{\prime}_{+}(v;1)$. This is in turn equivalent, by (3.20), to the fact that
$\min_{\xi\in\partial{\mathcal{R}}(v)}{\mathcal{R}}^{*}(\xi)=\max_{\xi\in\partial{\mathcal{R}}(v)}{\mathcal{R}}^{*}(\xi)\,,$
i.e. (3.16).
@@
with $\psi:[0,\infty)\to[0,\infty)$ convex with superlinear growth at infinity, as in Definition 2.1. For the associated mapping $f(v;\cdot)$, $v\in X$, we have
@@
Therefore, ${\mathcal{R}}$ is radially differentiable in the sense of (3.16) if and only if $\partial\psi$ is single-valued. We thus retrieve the smoothness requirement on $\psi$ from Hyp. 2.5.
Example 3.2 revisited.
Let us get back to the dissipation potential ${\mathcal{R}}:{\mathrm{L}}^{p}(\Omega)\to[0,\infty)$, $p\geq 1$, from Example 3.2, i.e. ${\mathcal{R}}(v)=\int_{\Omega}\mathrm{R}(v(x))\;\!\mathrm{d}x$, with the dissipation density $\mathrm{R}:{\mathbb{R}}\to[0,\infty)$ satisfying conditions (3.7). In that setting, for a given $v\in\mathrm{dom}(\partial{\mathcal{R}})$ we have that
@@
 is radially differentiable R is differentiable in {0}. ${\mathcal{R}}\text{ is radially differentiable }\ \Longleftrightarrow\ \mathrm{R}\text{ is differentiable in }{\mathbb{R}}{\setminus}\{0\}.$ (3.22)
Indeed, suppose that $\mathrm{R}:{\mathbb{R}}\to[0,\infty)$ is radially differentiable.
Then, in view of the characterization provided by Prop. 3.8, for all $v\in\mathrm{dom}(\partial{\mathcal{R}})$ and $\xi_{1},\,\xi_{2}\in\partial{\mathcal{R}}(v)$ we have
@@
The converse implication holds for, if $\mathrm{R}$ is not radially differentiable, then there exist $\hat{v},\hat{\xi}_{1},\hat{\xi}_{2}\in{\mathbb{R}}$ such that $\hat{\xi}_{i}\in\partial\mathrm{R}(\hat{v})$, $i=1,2$, and $\mathrm{R}^{*}(\hat{\xi}_{1})\neq\mathrm{R}^{*}(\hat{\xi}_{2})$. Then, defining $v(x)\equiv\hat{v}$ and $\xi_{i}(x)\equiv\hat{\xi}_{i}$ for almost all $x\in\Omega$, we obtain a triple $(v,\xi_{1},\xi_{2})\in{\mathrm{L}}^{p}(\Omega){\times}{\mathrm{L}}^{p^{\prime}}(\Omega){\times}{\mathrm{L}}^{p^{\prime}}(\Omega)$ for which (3.17) fails to hold.
@@
Ultimately, as a straightforward consequence of Prop. 3.8  we have the following result.
@@
Suppose that ${\mathcal{R}}:X\to[0,\infty)$ is radially differentiable. Then, for every $u\in J_{\sigma}(u^{o})$ we have that
$\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;u)={\mathcal{R}}^{*}({-}\xi)\qquad\text{for all }\,\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})\,.$
In particular, as soon as a (measurable) selection $\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ fulfills De Giorgi’s estimate/identity, then (3.13)/(3.14) hold with the slope part of the time integrated dissipation, namely $\int_{0}^{\sigma}\mathcal{C}_{\mathcal{R}}(u^{o},\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho$, given by $\int_{0}^{\sigma}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho$ for any (measurable) selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})$. Furthermore, if De Giorgi’s identity holds along an interval $(0,\sigma)$, the corresponding optimal integrand $\mathcal{C}_{\mathcal{R}}(u^{o},\rho;\widetilde{u}_{\rho})$ fulfills
$\mathcal{C}_{\mathcal{R}}(u^{o},\rho;\widetilde{u}_{\rho})=\frac{{\mathrm{d}}}{{\mathrm{d}}s}\big{(}s{\mathcal{R}}\big{(}\frac{1}{s}(\widetilde{u}_{\rho}{-}u^{0})\big{)}\big{)}|_{s=\rho}\qquad\text{for a.a.}\,\rho\in(0,\sigma)\,.$ (3.24)
In Theorem 4.1 ahead we will indeed prove De Giorgi’s identity under the condition that the dissipation potential is radially differentiable.
@@ 3.4 Tools
For fixed $v\in X$ define the function $g(t)=t{\mathcal{R}}\big{(}\frac{1}{t}\,v\big{)}$. Then, $g$ is convex, decreasing, and satisfies $g\in{\mathrm{C}}^{\mathrm{Lip}}_{\mathrm{loc}}({]0,\infty[})$. For all $t>0$ the left derivative $g^{\prime}_{-}(t)$ and the right derivative $g^{\prime}_{+}(t)$ exist, are non-decreasing, and satisfy
@@
For fixed $u^{o}\in X$, we consider the marginal function $(0,\infty)\ni\sigma\mapsto\phi(u^{o};\sigma)$ of $(\sigma,u)\mapsto\Phi_{\sigma}(u^{o};u)$. The point is that one-sided differentiability of $\Phi$ with respect to $\sigma$ provides bounds on the one-sided derivatives of $\phi$, for which the behavior of $\Phi$ with respect to $u$ is not really important. In the following result, in fact, we do not use the special form of $\Phi$, but only its left and right differentiability with respect to $\sigma$.
@@
We have $\phi\in{\mathrm{C}}^{\mathrm{Lip}}_{\mathrm{loc}}({]0,\infty[})$ and for all $\sigma>0$ the following estimates hold:
$\liminf_{h\to 0^{+}}\frac{1}{h}\big{(}\phi(u^{o};\sigma)-\phi(u^{o};\sigma{-}h)\big{)}\geq\sup\big{\{}\,{\mathrm{D}}^{-}_{\sigma}\Phi_{\sigma}(u^{o};w)\>\big{|}\>w\in J_{\sigma}(u^{o})\,\big{\}}=:\delta^{-}_{\sigma}(u^{o})$ (3.25a)
$\limsup_{h\to 0^{+}}\frac{1}{h}\big{(}\phi(u^{o};\sigma{+}h)-\phi(u^{o};\sigma)\big{)}\leq\inf\big{\{}\,{\mathrm{D}}^{+}_{\sigma}\Phi_{\sigma}(u^{o};w)\>\big{|}\>w\in J_{\sigma}(u^{o})\,\big{\}}=:\delta^{+}_{\sigma}(u^{o}).$ (3.25b)
Proof. For notational simplicity we drop the argument $u^{o}$ in $\phi$ and $\Phi$.
Consider $0<r<\sigma$, then by the marginal property of $\phi$ we have
@@
Hence, the local Lipschitz property of $\phi$ follows from that of $\Phi(\cdot,w)$. Setting $r=t{-}h$ yields
@@
Taking the liminf for $h\to 0^{+}$ first and then the supremum over $\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ gives (3.25a).
@@
Taking the limsup $h\to 0^{+}$ first and then the infimum over $\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ yields (3.25b).

### v3

@@ 3.2 The De Giorgi estimate and identity in the Banach setup
$\int_{0}^{\sigma}{\mathcal{R}}^{*}\big({-}\widetilde{\xi}_{\rho}\big)\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ})-{\mathcal{E}}(\widetilde{u}_{\sigma})-\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)\quad\text{with }\widetilde{\xi}_{\rho}\in\partial{\mathcal{E}}(\widetilde{u}_{\rho})$ (3.18)
holds for $\sigma\in(0,1]$, while for $\sigma>1$ it only holds for selections $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ with
$\int_{0}^{\sigma}{\mathcal{R}}^{*}\big({-}\widetilde{\xi}_{\rho}\big)\;\!\mathrm{d}\rho=\int_{0}^{1}\frac{1}{2}\;\!\mathrm{d}\rho+\int_{1}^{\sigma}\frac{1}{2}\big|{-}\widetilde{\xi}_{\rho}\big|^{2}\;\!\mathrm{d}\rho\leq 1-\frac{1}{2\sigma}\quad\Leftrightarrow\quad\int_{1}^{\sigma}\big|\widetilde{\xi}_{\rho}\big|^{2}\;\!\mathrm{d}\rho\leq 1-\frac{1}{\sigma}.$
Since $\mathcal{S}_{{\mathcal{R}}}(\widetilde{u}_{\rho})=0$ for $\rho\geq 1$, (3.18) is certainly satisfied if we replace the integrand ${\mathcal{R}}^{*}\big({-}\widetilde{\xi}_{s}\big)$ by $\mathcal{S}_{{\mathcal{R}}}(\widetilde{u}_{s})$; thus the simple De Giorgi estimate (3.15) is clearly satisfied.
However, recalling that for $\sigma\geq 1$ we have $\partial{\mathcal{E}}(\widetilde{u}_{\sigma})=[0,1]$, a wrong choice of $\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ can violate (3.18).
@@
$0\in\widetilde{\xi}_{\sigma}+\partial{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)\quad\text{with }\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma}).$
@@
$\widetilde{\xi}_{\sigma}=-\frac{1}{\sigma}(\widetilde{u}_{\sigma}-u^{\circ})=\begin{cases}1&\text{for }\sigma\in(0,1]\,,\\ 1/\sigma&\text{for }\sigma\geq 1\,.\end{cases}$
@@
$\text{for }\sigma\in(0,1]:$ $\quad\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})=\{1\}\quad\text{and}\quad{\mathcal{S}}_{\mathcal{R}}(\widetilde{u}_{\sigma})={\mathcal{C}}_{\mathcal{R}}(\sigma;\widetilde{u}_{\sigma})=\frac{1}{2},$
$\text{for }\sigma>1:$ $\quad\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})=\big\{\frac{1}{\sigma}\big\}\quad\text{and}\quad 0={\mathcal{S}}_{\mathcal{R}}(\widetilde{u}_{\sigma})\lneqq{\mathcal{C}}_{\mathcal{R}}(\sigma;\widetilde{u}_{\sigma})=\frac{1}{2\sigma^{2}}\>.$
@@
$\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})\qquad\text{for a.a.}\ \sigma>0\,.$
Thus, the ‘good selections’ of $\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ for the purpose of De Giorgi identity is prescribed by the Euler-Lagrange equation (3.7).
In summary, we have the simple De Giorgi estimate (3.15) for $\mathcal{S}_{{\mathcal{R}}}$ (with strict inequality for $\sigma>1$), and we have the De Giorgi identity (3.17) for $\mathcal{C}_{{\mathcal{R}}}$.
@@ 3.3 Radially differentiable potentials
This section revolves around a structural property for dissipation potentials that will be crucial for obtaining the De Giorgi identity.
@@
We say that a dissipation potential ${\mathcal{R}}:X\to[0,\infty]$ is radially differentiable if
$\forall\,v\in\mathop{\mathrm{dom}}({\mathcal{R}}):\quad(0,1)\ni\lambda\mapsto f(v;\lambda):={\mathcal{R}}(\lambda v)\text{ is differentiable. }$ (3.19)
Observe that, if $\mathrm{dom}({\mathcal{R}})$ is open, then (3.19) in fact implies that $\lambda\mapsto f(v;\lambda)$ is well-defined and differentiable on the interval $(0,1+\epsilon)$ for some $\epsilon>0$.
@@
We are now going to show that radial differentiability is equivalent to another structural property that was used in [MRS13] to prove the De Giorgi estimate (3.16).
@@
This ensures that the (a priori multivalued) mapping (whose domain is clearly given by $\mathop{\mathrm{dom}}(\partial{\mathcal{R}})$)
$\mathfrak{P}:X\rightrightarrows[0,\infty),\qquad\mathfrak{P}(v):=\langle\xi,v\rangle\text{ for all }\xi\in\partial{\mathcal{R}}(v),\quad\text{ is single-valued}$ (3.21)
@@
Let us now address the proof of the equivalence of (3.20) and (3.19): We start by observing that for every $v\in\mathop{\mathrm{dom}}({\mathcal{R}})$ the mapping $f(v;\cdot)$ is convex. Its subdifferential in the sense of convex analysis $\partial f(v;\cdot):(0,\infty)\rightrightarrows{\mathbb{R}}$ is given by
$\partial f(v;\lambda)=[f^{\prime}_{-}(v;\lambda),f^{\prime}_{+}(v;\lambda)]\qquad\text{for all }\lambda>0,\ v\in\mathop{\mathrm{dom}}({\mathcal{R}})\,$, (3.22a)
@@
$f^{\prime}_{+}(v;\lambda)=\lim_{h\to 0^{+}}\frac{1}{h}\big(f(v;\lambda{+}h)-f(v;\lambda)\big),$ (3.22b)
$f^{\prime}_{-}(v;\lambda)=\lim_{h\to 0^{+}}\frac{1}{h}\big(f(v;\lambda)-f(v;\lambda{-}h)\big)\,$,
(the above limits exist by convexity of $f(v;\cdot)$).
@@
For every $v\in\mathop{\mathrm{dom}}({\mathcal{R}})$ and $\lambda>0$
@@
Proof. It can be easily checked that, for any $v\in\mathop{\mathrm{dom}}({\mathcal{R}})$ and $\lambda>0$, there holds
@@
Proof of Proposition 3.11. The mapping $f(v;\cdot)$ from (3.19) is differentiable at $\lambda=1$ if and only if $f^{\prime}_{-}(v;1)=f^{\prime}_{+}(v;1)$. This is in turn equivalent, by (3.23), to the fact that $\min_{\xi\in\partial{\mathcal{R}}(v)}{\mathcal{R}}^{*}(\xi)=\max_{\xi\in\partial{\mathcal{R}}(v)}{\mathcal{R}}^{*}(\xi)$, i.e. (3.19).
@@
with $\psi:[0,\infty)\to[0,\infty)$ convex with superlinear growth at infinity, as in Definition 2.1. For the associated mapping $f(v;\cdot)$, $v\in X$ (note that $\mathop{\mathrm{dom}}({\mathcal{R}})=X$, in this case), we have
@@
Therefore, $\mathcal{R}$} is radially differentiable in the sense of (3.19) if and only if $\partial\psi$ is single-valued. We thus retrieve the smoothness requirement on $\psi$ from Hypothesis 2.5.
Remark 3.14 (Example 3.5 revisited. )
Let us get back to the dissipation potential ${\mathcal{R}}:{\mathrm{L}}^{p}(\Omega)\to[0,\infty)$, $p\geq 1$, from Example 3.5, i.e. ${\mathcal{R}}(v)=\int_{\Omega}\mathrm{R}(v(x))\;\!\mathrm{d}x$, with the dissipation density $\mathrm{R}:{\mathbb{R}}\to[0,\infty)$ satisfying conditions (3.8) (so that $\mathop{\mathrm{dom}}({\mathcal{R}})={\mathrm{L}}^{p}(\Omega)$). In that setting, for a given $v\in\mathop{\mathrm{dom}}(\partial{\mathcal{R}})$ we have that
@@
 ${\mathcal{R}}\text{ is radially differentiable }\ \Longleftrightarrow\ \mathrm{R}\text{ is differentiable in }{\mathbb{R}}{\setminus}\{0\}.$ (3.25)
Indeed, suppose that $\mathrm{R}:{\mathbb{R}}\to[0,\infty)$ is radially differentiable. Then, in view of the characterization provided by Prop. 3.11, for all $v\in\mathop{\mathrm{dom}}(\partial{\mathcal{R}})$ and $\xi_{1},\,\xi_{2}\in\partial{\mathcal{R}}(v)$ we have
@@
[…]
@@
Ultimately, as a straightforward consequence of Prop. 3.11 we have the following result.
@@
Suppose that ${\mathcal{R}}:X\to[0,\infty)$ is radially differentiable. Then, we have that
$\mathcal{C}_{{\mathcal{R}}}(\sigma;u)={\mathcal{R}}^{*}({-}\xi)\qquad\text{for all }\,\xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})\text{ and }u\in J_{\sigma}\,.$
In particular, as soon as a (measurable) selection $\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ fulfills the De Giorgi estimate/identity, then (3.16)/(3.17) hold with the slope part of the time integrated dissipation, namely $\int_{0}^{\sigma}\mathcal{C}_{{\mathcal{R}}}(\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho$, given by $\int_{0}^{\sigma}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho$ for any (measurable) selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})$. Furthermore, if the De Giorgi identity holds along an interval $(0,\sigma)$, the corresponding optimal integrand $\mathcal{C}_{{\mathcal{R}}}(\rho;\widetilde{u}_{\rho})$ fulfills
$\mathcal{C}_{{\mathcal{R}}}(\rho;\widetilde{u}_{\rho})=\frac{{\mathrm{d}}}{{\mathrm{d}}s}\big(s{\mathcal{R}}\big(\frac{1}{s}(\widetilde{u}_{\rho}{-}u^{\circ})\big)\big)|_{s=\rho}\qquad\text{for a.a.}\,\rho\in(0,\sigma)\,.$ (3.27)
In Theorem 4.2 ahead we will indeed prove the De Giorgi identity under the condition that the dissipation potential is radially differentiable.
@@ 3.4 Tools
For fixed $v\in\mathrm{dom}({\mathcal{R}})$ define the function $g(t)=t{\mathcal{R}}\big(\frac{1}{t}\,v\big)$. Then, $g$ is convex, decreasing, and satisfies $g\in{\mathrm{C}}^{\mathrm{Lip}}_{\mathrm{loc}}({]0,\infty[})$. For all $t>0$ the left derivative $g^{\prime}_{-}(t)$ and the right derivative $g^{\prime}_{+}(t)$ exist, are non-decreasing, and satisfy
@@
Recall that $(0,\infty)\ni\sigma\mapsto\phi(\sigma)$ is the marginal function of $(\sigma,u)\mapsto\Phi_{\sigma}(u)$. Our next result shows that one-sided differentiability of $\Phi$ with respect to $\sigma$ provides bounds on the one-sided derivatives of $\phi$, for which the behavior of $\Phi$ with respect to $u$ is not really important. In the following result, in fact, we do not use the special form of $\Phi$, but only its left and right differentiability with respect to $\sigma$.
@@
We have $\phi\in{\mathrm{C}}^{\mathrm{Lip}}_{\mathrm{loc}}({]0,\infty[})$ and for all $\sigma>0$ the following estimates hold
$\liminf_{h\to 0^{+}}\frac{1}{h}\big(\phi(\sigma)-\phi(\sigma{-}h)\big)\geq\sup\big\{\,{\mathrm{D}}^{-}_{\sigma}\Phi_{\sigma}(w)\>\big|\>w\in J_{\sigma}\,\big\}=:\delta^{-}_{\sigma},$ (3.28a)
$\limsup_{h\to 0^{+}}\frac{1}{h}\big(\phi(\sigma{+}h)-\phi(\sigma)\big)\leq\inf\big\{\,{\mathrm{D}}^{+}_{\sigma}\Phi_{\sigma}(w)\>\big|\>w\in J_{\sigma}\,\big\}=:\delta^{+}_{\sigma}$ (3.28b)
with ${\mathrm{D}}^{\pm}_{\sigma}\Phi_{\sigma}(w)$ the one-sided derivatives of the map $\sigma\mapsto\Phi_{\sigma}(w)$.
Proof. Consider $0<r<\sigma$, then by the marginal property of $\phi$ we have
@@
Hence, the local Lipschitz property of $\phi$ follows from that of $\Phi(\cdot,w)$. Setting r=thr=t{-}h yields
@@
Taking the liminf for $h\to 0^{+}$ first and then the supremum over $\widetilde{u}_{\sigma}\in J_{\sigma}$ gives (3.28a).
@@
Taking the limsup $h\to 0^{+}$ first and then the infimum over $\widetilde{u}_{\sigma}\in J_{\sigma}$ yields (3.28b).